If u = tan inverse ((x cube - y cube) /(x cube + y cube)) .Then show that x.ux + y.uy = 0
1
Expert's answer
2020-09-06T17:25:32-0400

As per the question,

u="tan^{-1}\frac{x^3-y^3}{x^3+y^3}"

differentiate u with respect to x

"\frac{du}{dx}=\frac{1}{1+(\frac{x^3-y^3}{x^3+y^3})^2}\times\frac{(x^3+y^3)3x^2-(x^3-y^3)3x^2}{(x^3+y^3)^2}"

on solving above expression we get,

"u_x=\frac{1}{(x^3+y^3)^2+(x^3-y^3)^2}\times(3x^2y^3+3x^2y^3)"

"u_x=\frac{1}{(x^3+y^3)^2+(x^3-y^3)^2}\times(6x^2y^3)" ....(1)

Now differentiating u with respect to y.

"\frac{du}{dy}=\frac{1}{1+(\frac{x^3-y^3}{x^3+y^3})^2}\times\frac{(x^3+y^3)(-3y^2)-(x^3-y^3)3y^2}{(x^3+y^3)^2}"

on solving we get,

"u_y=\frac{1}{(x^3+y^3)^2+(x^3-y^3)^2}\times(-3x^3y^2-3x^3y^2)"

"u_y=\frac{1}{(x^3+y^3)^2+(x^3-y^3)^2}\times(-6x^3y^2)" ...(2)


To show "xu_x+yu_y=0"

"\frac{x}{(x^3+y^3)^2+(x^3-y^3)^2}\times(6x^2y^3)" "+\frac{y}{(x^3+y^3)^2+(x^3-y^3)^2}\times(-6x^3y^2)"


="\frac{1}{(x^3+y^3)^2+(x^3-y^3)^2}\times(6x^3y^3-6x^3y^3)"

=0

Hence "xu_x+yu_y=0" .